### v1

# Bosch Research North America & Bosch Center for AI https://zixunh.github.io/3d-geer

Abstract.
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) marks a significant milestone in balancing the quality and efficiency of differentiable rendering. However, its high efficiency stems from an approximation of projecting 3D Gaussians onto the image plane as 2D Gaussians, which inherently limits rendering quality—particularly under large Field-of-View (FoV) camera inputs. While several recent works have extended 3DGS to mitigate these approximation errors, none have successfully achieved both exactness and high efficiency simultaneously. In this work, we introduce 3DGEER , an Exact and Efficient Volumetric Gaussian Rendering method. Starting from first principles, we derive a closed-form expression for the density integral along a ray traversing a 3D Gaussian distribution. This formulation enables precise forward rendering with arbitrary camera models and supports gradient-based optimization of 3D Gaussian parameters. To ensure both exactness and real-time performance, we propose an efficient method for computing a tight Particle Bounding Frustum (PBF) for each 3D Gaussian, enabling accurate and efficient ray-Gaussian association. We also introduce a novel Bipolar Equiangular Projection (BEAP) representation to accelerate ray association under generic camera models. BEAP further provides a more uniform ray sampling strategy to apply supervision, which empirically improves reconstruction quality. Experiments on multiple pinhole and fisheye datasets show that our method consistently outperforms prior methods, establishing a new state-of-the-art in real-time neural rendering.
Volumetric Rendering, Differentiable Rendering, Novel View Synthesis, Radiance Fields, Neural Reconstruction
journal: TOGccs: Computing methodologies Renderingccs: Computing methodologies Reconstructionccs: Computing methodologies Machine Learning Approaches22footnotetext: Corresponding Author.
1. Introduction
Initiated by 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023), volumetric particle rendering methods formulated under splatting (Yu et al., 2024a; Yan et al., 2024; Lu et al., 2024; Kheradmand et al., 2024; Charatan et al., 2024; Mallick et al., 2024) have gained significant popularity due to their impressive visual fidelity and high rendering speed. The core efficiency stems from EWA splatting (Zwicker et al., 2001), which uses a local affine approximation to project a 3D Gaussian as a 2D Gaussian on the image plane. While effective in narrow field-of-view (FoV) scenarios, this approximation introduces substantial errors in wide-FoV settings, where nonlinear distortions cause the true projection to deviate significantly from a symmetric 2D Gaussian—especially in fisheye images commonly encountered in robotics and autonomous driving (Rashed et al., 2021; Courbon et al., 2007; Kumar et al., 2023; Zhang et al., 2016; Yogamani et al., 2019, 2024; Sekkat et al., 2022). Although recent works have attempted to enable differentiable rendering under specific fisheye camera models (Liao et al., 2024) or more general camera models (Huang et al., 2024), these approaches remain fundamentally constrained by their reliance on 2D projective-space approximations and have yet to yield an exact closed-form solution.
A parallel line of work tackles volumetric particle rendering directly in a ray-marching formulation (Yu et al., 2024b; Condor et al., 2025; Mai et al., 2024; Moënne-Loccoz et al., 2024). These approaches naturally support large-FoV camera models and are free from projective approximation error. However, achieving high efficiency remains a key challenge—particularly in associating each rendering ray with its most relevant 3D particles. To address this, significant effort has been devoted to accelerating it, such as adopting advanced Bounding Volume Hierarchies (BVH) employed in EVER (Mai et al., 2024) and 3DGRT (Moënne-Loccoz et al., 2024). Nevertheless, due to the inherent algorithmic complexity and difficulty in parallelization, pure ray marching methods have yet to match the efficiency of 3DGS. Alternatively, a hybrid approach has been proposed in 3DGUT (Wu et al., 2025) which adopts ray-tracing rendering with rasterization to solve ray-particle association efficiently. However, the inherited projective approximation error from association stage still compromises rendering exactness. To date, no existing method achieves both the exactness of ray-based rendering and the high efficiency of projection-based splatting.
In this work, we propose 3DGEER (3D Gaussian Exact and Efficient Rendering)—a novel volumetric Gaussian rendering method that achieves exact rendering while maintaining high efficiency comparable to 3DGS (see Fig. 1). 3DGEER is built upon three key technical contributions: (i) Instead of relying on projective approximations or 2D conic fitting, we revisit ray-marching-based volumetric Gaussian rendering from first principles and derive a closed-form solution for integrating density along a 3D ray through a Gaussian. Based on it, we further derive the exact volumetric Gaussian rendering framework. (ii) To perform efficient ray-particle association without compromising rendering exactness, we reformulate the problem as associating Camera Sub-frustums (CSF) with Particle Bounding Frustums (PBF)—analogous to the tile-AABB mapping used in 3DGS. For each 3D Gaussian, we derive a closed-form solution to compute the PBF tightly bounding it, enabling highly efficient association. (iii) In addition, we propose to uniformly sample rays in a novel Bipolar Equiangular Projection (BEAP) space to apply color supervision. This ray sampling strategy not only aligns image-space partitioning and the underlying CSF but also improved conventional pinhole or equidistant projections in reconstruction quality. Extensive experiments on both fisheye and pinhole datasets—including ScanNet++ (Yeshwanth et al., 2023), Zip-NeRF (Barron et al., 2023), and MipNeRF (Barron et al., 2021)—show that 3DGEER consistently outperforms prior and concurrent methods, achieving 0.9–4.8 PSNR gains under challenging wide-FoV conditions, and 0.3–2.0 on pinhole views.
In summary, our contributions are as follows:
  1. (1)
    We present the first complete, first-principle derived, closed-form solution for exact volumetric Gaussian rendering;
  2. (2)
    We propose an exact and efficient ray-particle association method that enables high speed rendering;
  3. (3)
    We introduce an equiangular ray sampling strategy that provides more spatially uniform color supervision.
  4. (4)
    Our method consistently outperforms SOTA approaches across multiple datasets, with notable gains on wide-FoV neural reconstruction benchmarks.
Table 1. Summary of Volumetric Particle Rendering Methods wrt. Exactness. The top section lists splatting-based methods, while the bottom section presents ray-marching-based approaches—some of which still rely on projective approximations (e.g., EWA or Unscented Transform) during ray-particle association.
Method Render Particle Asso. Particle Large High Exact
Meth. Scope Type FoV FPS Meth.
3DGS (2023) Splat. EWA Img. 3D Gauss.
FisheyeGS (2024) Splat. EWA Img. 3D Gauss.
GS++  (2024) Splat. EWA Img. 3D Gauss.
GOF (2024b) Ray EWA Img. 3D Gauss.
EVER (2024) Ray BVH Scene 3D Ellip.
3DGRT (2024) Ray BVH Scene 3D Gauss.
3DGUT (2025) Ray UT Img. 3D Gauss.
3DGEER (Ours) Ray PBF Frust. 3D Gauss.
2. Related Works
2.1. Splatting-Based Gaussian Rendering
Most 3DGS variants  (Yu et al., 2024a; Yan et al., 2024; Lu et al., 2024; Kheradmand et al., 2024; Zhang et al., 2024; Rota Bulò et al., 2024; Chen et al., 2024b; Lee et al., 2024a; Kerbl et al., 2024; Guédon and Lepetit, 2024; Chen et al., 2024a; Chung et al., 2024; Lee et al., 2024b; Charatan et al., 2024; Mallick et al., 2024; Cheng et al., 2024) adopt the original EWA splatting (Zwicker et al., 2001) strategy, projecting 3D Gaussians onto the image plane and approximating their appearance as 2D conics. To support differentiable rendering under large-FoV camera models with distortion, methods such as FisheyeGS (Liao et al., 2024) and GS++ (Huang et al., 2024) estimate the Jacobian of equidistant or spherical projections to adjust the 2D mean and covariance of the projected Gaussians. However, these splatting-based approaches still rely on a first-order Taylor expansion of the highly non-linear projection function. In wide-FoV settings, higher-order terms is non-negligible, resulting in significant approximation errors and degraded reconstruction quality.
Fundamentally, the true projection of a 3D Gaussian under non-linear camera models—whether ideal pinhole or distorted projections—is not a symmetric 2D Gaussian. In other words, any framework that relies on linear projective geometry to approximate inherently nonlinear transformations will inevitably introduce approximation errors that cannot be fully eliminated. In contrast, our approach avoids such projective approximations entirely by aggregating Gaussian density directly along rays in a volumetric rendering formulation.
2.2. Ray-Marching-Based Gaussian Rendering
A few recent works have approached volumetric particle rendering using a ray-marching formulation, which inherently supports arbitrary camera models without introducing projective approximation error. 3DGRT (Moenne-Loccoz et al., 2024), inspired by fuzzy metaball rendering (Keselman and Hebert, 2022), approximates the ray-Gaussian interaction as the location of maximum intensity response. This heuristic enables ray-space sampling without explicit projection, effectively mitigating wide-FoV distortion. However, it incurs significant computational overhead due to the high cost of ray-particle association via scene-level spatial partitioning structures such as BVH, ultimately resulting in low frame rates.
Alternatively, EVER (Mai et al., 2024) replaces the Gaussian representation with simplified ellipsoids of constant density, enabling a closed-form solution for transmittance-weighted ray integration. However, it reduces the expressive capacity of the representation compared to full 3D Gaussians, and continues to suffer from low runtime efficiency due to its reliance on BVH-based structures.
Figure 2. Projective Approximation Error in Ray-Particle Association. When projecting a Gaussian particle from 3D space onto the image plane, both EWA and Unscented Transform (UT)-based association methods approximate the 2D covariance of the projection. They then either tighten the bounding box—risking mismatched culling—or overestimate it using the maximum projected radius of the conic. In contrast, our PBF method avoids this intermediate approximation by directly computing the exact bounding structure from the true 3D covariance.
3DGUT (Wu et al., 2025), a follow-up to 3DGRT, improves runtime efficiency by adopting splatting-based particle association, achieving frame rates comparable to 3DGS. However, its acceleration relies on the Unscented Transform (UT), a sampling-based approximation used to estimate 2D conics in projective image space. While the rendering stage is ray-based, the ray-particle association still introduces approximation errors akin to those in EWA splatting. A similar limitation exists in GOF (Yu et al., 2024b), which directly applies the EWA trick to fit 2D bounding conics and projective centers for association. Notably, under extreme FoV conditions, imprecise AABB estimation and the resulting inaccurate tile-to-AABB associations can lead to visible rendering artifacts (see Figure 2).
In contrast to these methods, our approach solves the ray-particle association problem directly at the frustum level in 3D space, preserving the exactness of ray-based rendering across arbitrary camera models while maintaining rendering speeds comparable to 3DGS. Table 1 summarizes representative methods that aim to address projective approximation errors, highlighting key differences across several important dimensions.
3. 3DGEER
Our method, 3DGEER, is built upon three key components: an exact ray-based volumetric Gaussian rendering formulation, an efficient frustum-particle association strategy, and an effective image representation. In Sec.3.1, we introduce the necessary preliminaries. Sec.3.2 presents the closed-form solution for Gaussian density integral along a 3D ray, followed by our differentiable volumetric rendering method that supports arbitrary FoV camera models. In Sec.3.3, we propose a novel and efficient approach for ray-particle association based on fast, exact computation of the Particle Bounding Frustum (PBF). Finally, in Sec.3.4, we detail the Bipolar Equiangular Projection (BEAP) image representation, which seamlessly supports generic camera models and accelerates frustum-particle association while improving reconstruction quality in practice.
3.1. Preliminaries
We follow the classic parameterization of 3D Gaussian primitives as introduced in prior work (Zwicker et al., 2001). Each 3D Gaussian distribution is defined by a mean position $\boldsymbol{\mu}\in\mathbb{R}^{3}$ and a 3D covariance matrix $\Sigma$ in the world coordinate system:
(1) $\mathcal{G}_{\Sigma,\boldsymbol{\mu}}(\mathbf{x})=\frac{1}{\rho\left|\Sigma\right|^{1/2}}\exp\left(-\frac{1}{2}(\mathbf{x}-\boldsymbol{\mu})^{\top}\Sigma^{-1}(\mathbf{x}-\boldsymbol{\mu})\right),$
where we adopt the normalization constant $\rho$ as $\sqrt{2\pi}$, and $\mathbf{x}\in\mathbb{R}^{3}$ denotes a point in world space. This formulation differs from recent methods such as 3DGS (Kerbl et al., 2023) and 3DGUT (Wu et al., 2025), which omit the determinant term $|\Sigma|^{1/2}$ to simplify splatting or association during rendering. The form of our covariance matrix obeys: $\Sigma=RSS^{\top}R^{\top}$ , given a scaling matrix $S=\mathrm{diag}(\mathbf{s})$ constructed from the scaling vector $\mathbf{s}\in\mathbb{R}^{3}$, and a rotation matrix $R\in\mathrm{SO}(3)$ derived from a unit quaternion $\mathbf{q}\in\mathbb{R}^{4}$.
To evaluate the transmittance contributed by a single 3D Gaussian particle along a ray $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$, we have:
[…]
where $\sigma$ is the opacity coefficient associated with the Gaussian.
Finally, considering a set of 3D Gaussian particles sorted by depth contributing to the rendering color of the given ray, the rendered color $C$ is computed as:
(3) $C(\mathbf{o},\mathbf{d})=\sum_{i}c_{i}(\mathbf{o},\boldsymbol{\mu}_{i})T_{i}\prod_{j=1}^{i-1}(1-T_{j}),$
where $c_{i}(\mathbf{o},\boldsymbol{\mu}_{i})$ derives the view-dependent color of the $i$-th Gaussian as seen from the optical center $\mathbf{o}$, parameterized using spherical harmonics (SH) (Kerbl et al., 2023). Through minimizing the reconstruction loss between the rendered color and the observed color along each ray, the parameters $\mathbf{s},\mathbf{q},\boldsymbol{\mu}$, the coefficient $\sigma$ as well as the SH coefficients, are jointly optimized via backpropagation.
3.2. Exact Volumetric Gaussian Rendering
The core challenge in achieving exact volumetric Gaussian rendering in a ray marching formulation lies in obtaining a closed-form solution for integrating Gaussian density and color along a 3D ray. To address this, we first transform the ray—relative to each Gaussian—into a canonical coordinate system where the transmittance integral (see Eq. B.4) remains consistent and the Gaussians become isotropic. We then derive the exact ray color from first principles.
Specifically, we define a canonical coordinate system shared across all 3D Gaussian primitives. A point $\mathbf{u}\in\mathbb{R}^{3}$ in this canonical space is mapped to world coordinates for each Gaussian via:
(4) $\begin{bmatrix}\mathbf{x}\\ 1\end{bmatrix}$ $=H\begin{bmatrix}\mathbf{u}\\ 1\end{bmatrix}=\begin{bmatrix}RS&\boldsymbol{\mu}\\ \mathbf{0}&1\end{bmatrix}\begin{bmatrix}\mathbf{u}\\ 1\end{bmatrix},$
Figure 3. Integral of 3D Gaussian Density along a Ray. Under the transformation from the world frame to the canonical frame, the ray integral remains invariant. The green box highlights the product of density and ray segment length (see Eq. 5).
Figure 4. Pipeline Flowchart. Illustration of forward and backward gradient propagation from the Gaussian parameters $\mathbf{s}$, $\mathbf{q}$, $\boldsymbol{\mu}$, and the coefficient $\sigma$ to the transmittance.
From this transformation, an infinitesimal segment $\Delta\mathbf{u}$ along a ray in the canonical frame corresponds to a segment $\Delta\mathbf{x}=RS\Delta\mathbf{u}$ in the world frame.
To ensure that the transmittance contributed by each Gaussian along the ray is preserved across coordinate systems, we require that the elementary transmittance (i.e., the product of density and segment length) remains invariant under the transformation (see Figure 3):
(5) $\mathcal{G}_{\mathbf{I},\mathbf{0}}(\mathbf{u})\,\left|\Delta\mathbf{u}\right|=\mathcal{G}_{\Sigma,\boldsymbol{\mu}}(\mathbf{x})\,\left|\Delta\mathbf{x}\right|.$
This constraint implies that the Jacobian determinant is absorbed into the measure. Note that $\left|\frac{\Delta\mathbf{x}}{\Delta\mathbf{u}}\right|=\left|RS\right|=\left|\Sigma\right|^{1/2}$. We can thus express the corresponding canonical density as:
(6) $\mathcal{G}_{\mathbf{I},\mathbf{0}}(\mathbf{u})=\frac{1}{\rho}\exp\left(-\frac{1}{2}\mathbf{u}^{\top}\mathbf{u}\right),$
which is isotropic and shared across all primitives in canonical space.
This leads to an important conclusion that the exact closed-form solution for the transmittance integral of each Gaussian along the ray $\mathbf{r}(t)$ can thus be obtained by computing the accumulated density along the transformed ray $\mathbf{r}_{u}(t)=\mathbf{o}_{u}+t\mathbf{d}_{u}$ in the canonical isotropic Gaussian. This admits a closed-form solution in terms of the perpendicular distance from the ray to the Gaussian center, as captured in Theorem B.1. (see Appendix B for a complete proof):
[…]
where $\mathrm{D}_{\mathbf{\mu},\Sigma}(\mathbf{o},\mathbf{d})$ denotes the perpendicular distance from the Gaussian center (i.e., the origin in the canonical space) to the transformed ray $\mathbf{r}_{u}(t)$. Intuitively, this quantity also represents the minimal Mahalanobis distance from any points on the ray to the Gaussian in world space, measuring how close the ray passes by the particle distribution.
The remaining question is just about how to compute the perpendicular distance from the canonical origin to the transformed ray. We derive the transformed ray’s direction vector $\mathbf{d}_{u}$ and moment vector $\mathbf{m}_{u}$ as:
(8) $\mathbf{d}_{u}=\left(S^{-1}R^{\top}\right)\mathbf{d}\ ,\quad\mathbf{m}_{u}=\mathbf{o}_{u}\times\mathbf{d}_{u},$
where $\mathbf{o}_{u}=\left(S^{-1}R^{\top}\right)\left(\mathbf{o}-\boldsymbol{\mu}\right)$ denotes the optical center in canonical space. Based on this, the square of $\mathrm{D}_{\mathbf{\mu},\Sigma}(\mathbf{o},\mathbf{d})$ can be efficiently calculated as:
(9) $\left(\mathrm{D}_{\mathbf{\mu},\Sigma}(\mathbf{o},\mathbf{d})\right)^{2}=\frac{\mathbf{m}_{u}^{\top}\mathbf{m}_{u}}{\mathbf{d}_{u}^{\top}\mathbf{d}_{u}}.$
By substituting this distance square into Eq. 7, and then the derived transmittance value into Eq. 3, we compute the accumulated color along the ray.
To minimize the computational overhead caused by automatic differentiation during training, we explicitly derive the ray-integral function and the gradients from the transmittance to all the Gaussian parameters, specifically $\frac{\partial T}{\partial\mathbf{s}},\frac{\partial T}{\partial\mathbf{q}},\frac{\partial T}{\partial\boldsymbol{\mu}}$. The detailed derivations for these derivatives are provided in Appendix C and visually explained in the Figure 4.
3.3. Efficient Ray-Particle Association via PBF
Given an exact volumetric Gaussian rendering method, both its final accuracy—particularly under large FoV camera models—and its runtime efficiency depend heavily on the ray-particle association strategy. To determine which Gaussians contribute to each ray, we simplify the association problem by mapping Camera CSFs to PBFs—the tight frustums enclosing each Gaussian at a certain distance from its center. A Gaussian is assigned to a CSF if its PBF intersects the CSF. All Gaussians associated with a CSF are then considered contributors to every ray within that CSF.
Figure 5. An Illustration of PBF-CSF Association. Left: Each PBF and CSF is parameterized by four spherical angles, with the PBF additionally defined by four planes tangent to a 3D Gaussian ellipsoid. Right: The intersection between the PBF and CSF, unfolded onto the BEAP imaging plane.
To efficiently compute a PBF, we first define two spherical angles (See Fig. A.1 in Appendix A) in the camera space as:
(10) $\theta=\arctan\left(\frac{\mathbf{d}_{c,x}}{\mathbf{d}_{c,z}}\right),\quad\phi=\arctan\left(\frac{\mathbf{d}_{c,y}}{\mathbf{d}_{c,z}}\right),$
where $(\mathbf{d}_{c,x},\mathbf{d}_{c,y},\mathbf{d}_{c,z})$ denotes the normalized direction vector pointing to the camera (view) space point $\mathbf{d_{c}}$. The angles $\theta$ and $\phi$ correspond to the ray’s incidence angles in the horizontal ($xz$) and vertical ($yz$) planes, respectively, each ranging from $-\frac{\pi}{2}$ to $\frac{\pi}{2}$.
We define a PBF by four planes — two aligned with the $\theta$ angle and two with the $\phi$ angle (See Figure 5):
(11) $\mathbf{g}_{\theta_{1,2}}=\left(-1,0,\tan\theta_{1,2},0\right)^{\top},\quad\mathbf{g}_{\phi_{1,2}}=\left(0,-1,\tan\phi_{1,2},0\right)^{\top},$
where $\theta_{1}<\theta<\theta_{2}$ and $\phi_{1}<\phi<\phi_{2}$ specify the angular limits.
Given a camera with extrinsic matrix $V=\left[R_{c}\mid\mathbf{t}_{c}\right]\in\mathrm{SE}(3)$ and a canonical-to-world transformation matrix $H$ (See Eq. 4), we apply the inverse transpose of the camera-to-canonical matrix $(VH)^{-1}$ to parameterize the corresponding canonical planes using the angular limits defined in view space. This transformation enables us to effectively derive the constraints for the PBFs. The canonical planes are given by:
(12) $\mathbf{g}_{u_{1,2}}$ $=(VH)^{\top}\mathbf{g}_{\theta_{1,2}}=-(VH)_{0}+\tan\theta_{1,2}(VH)_{2},$
$\mathbf{g}_{v_{1,2}}$ $=(VH)^{\top}\mathbf{g}_{\phi_{1,2}}=-(VH)_{1}+\tan\phi_{1,2}(VH)_{2},$
where $(VH)_{i}$ denotes the $i$-th row of the matrix, expressed as a column vector.
Assuming a Gaussian ellipsoid defined by a $\lambda$-standard-deviation contour, we compute the angular bounds of the local frustum by solving the following quadratic equations, which enforce that the transformed planes are tangent to the ellipsoidal contour:
(13) $\left((\lambda^{2},\lambda^{2},\lambda^{2},-1)\circ\mathbf{g}_{u}^{\top}\right)\cdot\mathbf{g}_{u}=0,$
where $\circ$ denotes the Hadamard (element-wise) product, and $\lambda$ is a parameter that controls the spread of the contour, typically set to $3$. An identical constraint is applied to $\mathbf{g}_{v}^{\top}$ in order to compute the corresponding angular bounds $\phi_{1,2}$.
The roots of these quadratic equations determine the tangent values of the angular bounds $[\theta_{1},\theta_{2}]$ and $[\phi_{1},\phi_{2}]$, which define the PBF angular limits. Notably, solving these quadratic equations under canonical space constraints respects the true 3D covariance and 3D mean of the Gaussian. This formulation is both efficient and eliminates the need for any intermediate approximation steps as adopted in EWA or UT (See Appendix E.1 for details).
3.4. Bipolar Equiangular Projection (BEAP)
We introduce a novel BEAP image representation for several reasons. First, it effectively represents images from large field-of-view (FoV) cameras without introducing FoV loss. Second, it allows image tiles and their corresponding Camera Sub-frustums (CSFs) to share an identical parameterization in terms of two $\theta$ and two $\phi$ angles, enabling highly efficient PBF association and optimal GPU parallelization. Third, the resulting ray sampling achieves a more uniform distribution in 3D space compared to conventional projective rendering, which empirically improves training efficiency and reconstruction quality, as shown in the ablation study in Sec. 5.2.
Figure 6. Comparison of Ray Distributions under Varying Projections. Each point represents a sampled ray, defined by the intersection between a ray from the optical center and the sphere surface. Our BEAP projection yields the most uniform ray distribution across the sphere surface.
Specifically, we evenly sample the rays using the spherical angular coordinates ($\theta$, $\phi$) (See Eq. 10), and then map the rays back to a discrete $\mathrm{w}\times\mathrm{h}$ image space using a linear transformation consisting of a scaling factor and a center shift as:
(14) $(x,y)=\left\lfloor\frac{\mathrm{w}\theta}{\mathrm{FoV}_{x}}+\frac{\mathrm{w}+1}{2},\frac{\mathrm{h}\phi}{\mathrm{FoV}_{y}}+\frac{\mathrm{h}+1}{2}\right\rfloor,$
where $\mathrm{FoV}_{x},\mathrm{FoV}_{y}$ indicate the camera’s horizontal and vertical field-of-view. As illustrated in Figure 6, sampling rays uniformly in $(\theta,\phi)$ leads to more balanced spatial coverage within the view frustum. Our sampling strategy contrasts with projective sampling, which disproportionately allocates samples toward peripheral regions, and equidistant projection, which oversamples near the image center.
4. Experiments
Table 2. Quantitative Results on the ScanNet++ Dataset. Each method is trained using full FoV data but evaluated separately on the full FoV, central, and peripheral regions to assess performance under varying distortion levels. The region-wise metrics is weighted by pixel-wise masks, where LPIPS is omitted as it is not directly separable by region.
Method Full FoV Central Peripheral
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
FisheyeGS 27.81 0.946 0.139 32.44 0.956 23.28 0.914
EVER 29.47 0.924 0.167 29.93 0.924 28.72 0.925
3DGUT 30.64 0.944 0.150 31.87 0.945 28.84 0.937
3DGEER (Ours) 31.50 0.953 0.126 32.64 0.955 28.94 0.945
Table 3. Quantitative Results on the ZipNeRF Dataset. The experiments follow a cross-camera generalization setup, where each method is trained on either Fisheye (FE) or Pinhole (PH) data and evaluated separately on both Fisheye and Pinhole test sets.
Train Data $1/8$ FE $1/8$ FE $1/8$ FE $1/4$ PH $1/4$ PH $1/4$ PH
Test Data $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
FisheyeGS 23.18 26.44 0.858 0.868 0.211 0.239 19.43 26.61 0.791 0.889 0.247 0.212
EVER 25.17 26.14 0.880 0.851 0.153 0.237 22.90 26.45 0.835 0.862 0.207 0.222
3DGUT 24.77 25.59 0.879 0.804 0.183 0.324 18.61 25.96 0.662 0.841 0.300 0.270
3DGEER (Ours) 26.24 27.62 0.897 0.888 0.140 0.214 23.39 27.61 0.846 0.863 0.209 0.254
In this section, we evaluate our proposed method, 3DGEER, across multiple datasets featuring both pinhole and fisheye camera models.
Baselines. We compare 3DGEER with leading volumetric particle rendering methods targeting projective approximation errors in 3DGS (Kerbl et al., 2023). For large FoV fisheye inputs, we evaluate against FisheyeGS (Liao et al., 2024), EVER (Mai et al., 2024), and 3DGUT (Wu et al., 2025). Under pinhole settings, we additionally compare with 3DGS and 3DGRT (Moënne-Loccoz et al., 2024).
Datasets. To systematically evaluate performance under large FoV conditions, where projective approximation errors are most critical. We use five diverse scenes from the ScanNet++(Yeshwanth et al., 2023) dataset and the fisheye ZipNeRF(Barron et al., 2023), which has been largely overlooked in prior work. For conventional pinhole settings, we use the standard ZipNeRF and the widely adopted MipNeRF360 (Barron et al., 2021) datasets.
[…]
Implementation Details. To ensure fair comparison, we align our implementation with the original 3DGS in terms of initialization and hyperparameters on the ScanNet++ and MipNeRF360 datasets. For ZipNeRF, we adopt the training configurations of EVER and SMERF (Duckworth et al., 2024) for all candidate methods. Similar to most ray-based frameworks, our method does not directly access gradients of mean2D to control densification or pruning. Instead, we optimize using gradients of the view-space mean3D. Unless otherwise noted, all models are trained with consistent supervision strategies and architectural settings. For fisheye datasets (ScanNet++ and ZipNeRF), we precompute bijective grid mappings (Guo et al., 2025) and reproject all rendered outputs to the native KB-Fisheye image space (Kannala and Brandt, 2006) before evaluation. Experiments on ScanNet++ and MipNeRF360 are conducted using full-resolution images. For ZipNeRF, due to the high frame count and extremely large fisheye resolution, we use the maximum resolution (see Tab. 3) that allows all baselines to run reliably on a local desktop setup. Specifically, all experiments are conducted using an RTX 4090 GPU and 64GB of RAM as the default configuration.
Table 4. Quantitative Results on the MipNeRF360 Dataset. In addition to quality metrics, FPS is reported using an RTX 4090 as the default benchmark. Numbers sourced from original works are marked with (measured on RTX 6000 Ada) and (measured on RTX 5090).
[…]
3DGS 27.21 0.815 0.214 343
EVER 27.51 0.825 0.233 36
3DGRT 27.20 0.818 0.248 52 / 68
3DGUT 27.26 0.810 0.218 265 / 317
3DGEER (Ours) 27.76 0.821 0.210 327
4.1. Evaluation on Large-FoV Novel View Synthesis
We evaluate 3DGEER against representative baselines on two fisheye datasets designed for benchmarking novel view synthesis. To ensure a more comprehensive evaluation, we incorporate additional analysis tailored to each dataset.
ScanNet++. As shown in Tab. 2, our method consistently outperforms all prior methods across all metrics, with clear margins. To further analyze each method’s performance under varying distortion, we separately evaluate reconstruction quality in the central and peripheral image regions (Detailed rule explained in Appendix G). As expected, FisheyeGS—while effective at enabling differentiable rendering for fisheye inputs—does not explicitly address projective approximation error. It performs well in the central region but degrades significantly in the periphery. Although all methods show some performance drop in peripheral areas, 3DGEER achieves the best results in both regions. Qualitatively, as illustrated in Fig. 7, 3DGEER delivers superior peripheral reconstruction compared to FisheyeGS, exhibits fewer artifacts and sharper details than EVER, and produces crisper object boundaries and textures—such as wires, keyboards, and wooden floors—when compared to 3DGUT.
ZipNeRF. Since the ZipNeRF dataset provides frame-wise aligned pinhole and fisheye datasets, we go beyond standard evaluation on the fisheye dataset by also conducting a cross-camera generalization evaluation—training on fisheye views and evaluating on pinhole, and vice versa—to further assess each method’s robustness and adaptability. As shown in Tab. 3, 3DGEER consistently outperforms all other methods across all metrics when trained on fisheye data, regardless of whether the test set is fisheye or pinhole. This highlights 3DGEER’s ability to fully leverage the advantages of large FoV training data through exact volumetric rendering. When training on pinhole data, 3DGEER also achieves significant PSNR improvements on both fisheye and pinhole test sets. In contrast, FisheyeGS performs well when both training and testing are done on pinhole images, but its performance drops sharply when evaluated on fisheye inputs. Similarly, 3DGUT exhibits a substantial decline when tested on fisheye views after training on narrower FoV data. Overall, EVER and 3DGEER demonstrate the most robust generalization across camera models—attributed to their preservation of geometric exactness—while 3DGEER consistently outperforms EVER across all scenarios. Qualitatively, as shown in Fig. 9, 3DGEER reconstructs finer details and exhibits greater robustness to artifacts than competing methods, both in challenging peripheral regions and cross-camera evaluation settings.
4.2. Evaluation on Normal-FoV Novel View Synthesis
We further evaluate 3DGEER on the MipNeRF360 dataset to demonstrate its effectiveness under standard, narrow-FoV conditions. As shown in Tab. 4, 3DGEER achieves strong performance across all metrics, setting a new state-of-the-art. This evaluation also includes a runtime comparison in terms of FPS. Notably, 3DGEER is over 5$\times$ more efficient than other exact ray-marching methods such as EVER and 3DGRT, and is the only exact method with runtime comparable to the original 3DGS. Additionally, it consistently outperforms projective-approximation-based methods—including 3DGS and 3DGUT—in PSNR, SSIM, and LPIPS.
4.3. Comprehensive Runtime Analysis
Table 5. Runtime Analysis vs. Other Methods. 3DGEER is compared with efficient splatting-based methods on MipNeRF360 and ScanNet++, with a thorough comparison in stage-wise and total runtime.
Dataset Method Avg. Timings (ms) $\downarrow$
Prep. Dup. Sort Render Total
MipNeRF360 3DGS 0.32 0.27 0.65 1.40 2.92
3DGEER (Ours) 0.37 0.17 0.27 2.10 3.06
ScanNet++ FisheyeGS 0.10 0.63 1.45 2.33 4.70
3DGEER (Ours) 0.13 0.26 0.59 2.89 3.98
Table 6. Runtime Analysis: Asso. Method on ScanNet++. The efficiency of the PBF-based association is compared to the splatting-based EWA and UT with acceleration option SnugBox integrated into 3DGEER pipeline.
3DGEER Avg. Timings (ms) $\downarrow$ Avg.
Asso. Prep. Dup. Sort Render Total FPS $\uparrow$
EWA 0.13 1.17 3.27 12.69 17.84 56
+ SnugBox 0.13 0.60 1.42 6.22 8.68 115
UT 0.13 0.71 1.94 7.97 11.16 90
+ SnugBox 0.13 0.28 0.64 3.16 4.38 228
PBF (Ours) 0.13 0.26 0.59 2.89 3.98 251
We begin by analyzing runtime performance, comparing 3DGEER with two highly efficient splatting-based methods that share similar processing stages: 3DGS on the MipNeRF360 dataset and FisheyeGS on the ScanNet++ dataset. As shown in Tab. 6, 3DGEER incurs lower overhead in Gaussian duplication and sorting, offsetting the additional cost of ray-based rendering. As a result, it achieves overall runtime efficiency comparable to these optimized splatting methods.
Second, we evaluate the runtime of our ray-particle association strategy against projection-based alternatives—EWA (Zwicker et al., 2001) and UT (Wu et al., 2025)—within the 3DGEER framework. As shown in Tab. 6, our method consistently achieves a 2.5$\times$ to 5$\times$ speedup. This gain arises because projection-based schemes estimate screen-space bounds using first-order Taylor expansions or UT sampling, which often conservatively overestimate the extent using the maximum projected radius—leading to excessive overdraw and inefficient memory access in ray-based pipelines.
Even when accelerated with techniques like SnugBox (Hanson et al., 2025), projection-based methods remain fundamentally constrained by their reliance on approximated density centers. The resulting bounding regions are often misaligned with the true ray-projected distributions, leading to systematic inaccuracies (see Fig. 2). In fact, SnugBox can exacerbate these mismatches by tightening bounds around miscentered projections, further degrading visual fidelity, as shown in Fig. 8. Despite these refinements, such methods still fall short in speed compared to our approach.
5. Ablation Studies
Table 7. Impact of PBF-CSF Association. The evaluation employs the artifact-sensitive LPIPS metric. To assess cross-model robustness, each variant is also tested using models trained with the other association strategies. Additionally, we report the performance gap between in-method and cross-method evaluations to quantify consistency and generalization.
Asso. Method LPIPS $\downarrow$ Cross-LPIPS $\downarrow$ $\Delta$LPIPS(1e-2) $\downarrow$
EWA 0.1250 0.1321 0.71
UT 0.1249 0.1302 0.53
PBF (Ours) 0.1245 0.1251 0.06
5.1. Impact of Particle Bounding Frustum (PBF).
We evaluate the effectiveness of the proposed PBF-based ray-particle association strategy under challenging large-FoV distortion settings using the ScanNet++ dataset. To quantify perceptual degradation, we adopt LPIPS (Zhang et al., 2018) as our primary metric, given its high sensitivity to structural and high-frequency artifacts—issues often overlooked by PSNR and SSIM. We compare our PBF-based association to two representative alternatives: the EWA-style 2D conic projection on the perspective image plane (Zwicker et al., 2001), and the Unscented Transform (UT) approximation (Wu et al., 2025) in non-linear angular space. All methods are evaluated within the same 3DGEER training pipeline to ensure a modular and fair comparison.
As shown in Table 7, our PBF method achieves the lowest LPIPS, outperforming both projective and unscented approximations, which are more sensitive to projective distortion. In addition, we conduct cross-method evaluations by testing one method’s rendering using Gaussian fields trained under a different association scheme. This setup allows us to assess both cross-method robustness and method consistency. Our PBF method demonstrates significantly better generalization across these settings, indicating benefit of eliminating projective approximation errors at the association stage. These results also highlight 3DGEER’s potential as an universal rendering method that can improve novel view synthesis quality, even when applied to Gaussian fields reconstructed by other methods.
Qualitatively, Fig.8 (Right) shows both projective and unscented bounding strategies exhibit aliasing and visibility artifacts, particularly at object boundaries and occlusions. Notably, similar artifacts are observed in the original 3DGUT implementation (Wu et al., 2025), as illustrated in Fig. 8 (Left). In contrast, our frustum-based association produces clean and artifact-free renderings—even when used with Gaussian fields trained under alternative bounding schemes.
5.2. Impact of Bipolar Equiangular Projection (BEAP)
Table 8. Impact of BEAP. The ablation compares BEAP against the Perspective representation derived from undistorted fisheye images, focusing on the central region using mask-weighted evaluation metrics.
Method Supervison Central Region
Space PSNR $\uparrow$ SSIM $\uparrow$
3DGS Perspective 31.58 0.951
3DGEER w/o BEAP Perspective 32.51 0.954
3DGEER (Default) Angular 33.23 0.957
We evaluate the effectiveness of BEAP by comparing it against standard perspective image representations within our 3DGEER pipeline, as well as the original 3DGS pipeline. To introduce a controlled level of challenge from large-FoV scenarios, we conduct experiments on rectified and centrally cropped images from ScanNet++ scenes, where fisheye images are undistorted into wide-FoV pinhole views. For fair comparison, all rendered outputs are reprojected into the native ScanNet++ coordinate space, and pixels falling outside the rectified FoV are masked during PSNR and SSIM evaluation. As shown in Table 8, BEAP—by supervising rays uniformly sampled in angular space—leads to significantly improved reconstruction fidelity over perspective-based sampling.
6. Conclusions
We presented 3DGEER, the first complete, closed-form solution for exact volumetric Gaussian rendering derived from first principles. To complement this formulation, we introduced an efficient PBF computation method for ray-particle association, enabling high-speed rendering while maintaining exactness. Additionally, we proposed the BEAP image representation to facilitate effective color supervision under wide-FoV camera models. Extensive experiments demonstrate that 3DGEER consistently outperforms state-of-the-art methods across multiple datasets, with particularly strong results in challenging wide-FoV scenarios.
Appendix A Extended Visual Explanations
Figure A.1. An Illustration of the Two Spherical Angles Defining PBF and BEAP. A view-space direction $\mathbf{d}_{c}$ is first normalized onto the optical sphere (green spade), followed by angular projections onto the $xz$ and $yz$ planes. The orange point denotes the resulting coordinate $(\theta,\phi)$ on the unfolded BEAP imaging plane.
Appendix B Mathematical Proof to Ray-Gaussian Transmittance in Canonical Space
Theorem B.1 (Ray Integral of Isotropic 3D Gaussian).
Given a ray and a standard isotropic 3D Gaussian, the ray-Gaussian integral has a closed-form proportional to:
$\exp\left(-\frac{1}{2}\mathrm{D}^{2}\right),$
where $\mathrm{D}$ is the perpendicular distance from the Gaussian center to the ray.
Proof.
We first apply a translation to the coordinate system such that the Gaussian $\mathcal{G}$ is centered at the origin, and the ray $\mathbf{r}_{u}$ is translated accordingly. Next, we rotate the entire scene so that the ray direction is aligned with the $z$-axis. Due to the translation-invariance of the integral and the isotropy of the standard Gaussian, these transformations do not affect the result of the integral.
The translated standard isotropic 3D Gaussian is defined as:
(B.1) $\mathcal{G}_{\mathbf{I},\mathbf{0}}(\mathbf{u})=\frac{1}{\rho}\exp\left(-\frac{1}{2}\mathbf{u}^{\top}\mathbf{u}\right),$
which matches Eq. 6. Let the $z$-axis aligned ray as:
(B.2) $\mathbf{r}_{u}(t)=\begin{bmatrix}\mathbf{o}_{u}\\ \mathbf{o}_{v}\\ 0\end{bmatrix}+t\begin{bmatrix}0\\ 0\\ 1\end{bmatrix},$
the ray-Gaussian integral becomes:
[…]
$=\frac{\sqrt{2\pi}}{\rho}\exp\left(-\frac{1}{2}\mathrm{D}^{2}\right),$
where $\mathrm{D}^{2}=\frac{u_{0}^{2}+v_{0}^{2}}{2}$ denotes the squared perpendicular distance from the ray to the Gaussian center in canonical space. Thus, the integral is proportional to $\exp\left(-\frac{1}{2}\mathrm{D}^{2}\right)$, as stated in theorem. ∎
Note that the density function shared by our 3D primitives in the canonical space (see Eq. 6) follows the same formulation as in the isotropic case. Given the world-coordinate ray $(\mathbf{o},\mathbf{d})$ as the input of $\mathrm{D}$, its value also depends on the world-space Gaussian parameters $\boldsymbol{\mu},\Sigma$, which are used to transform the ray into the canonical coordinate system. Therefore, for each Gaussian contributing to the ray, we derive the transmittance as:
(B.4) $T=\frac{\sqrt{2\pi}\sigma}{\rho}\exp\left(-\frac{1}{2}\left(\mathrm{D}_{\boldsymbol{\mu},\Sigma}(\mathbf{o},\mathbf{d})\right)^{2}\right),$
which recovers Eq. 7 when $\rho=\sqrt{2\pi}$.
Appendix C Cookbook of Gradient Computation
C.1. Chain Rule
Recall that $T=\sigma\alpha$ is the ray-Gaussian transmittance (Eq. 7), where $\sigma$ is the opacity coefficient, and $\alpha$ is the ray-Gaussian integral. We denote $\kappa=\mathrm{D}^{2}_{\mu,\Sigma}$ as the minimal squared Mahalanobis distance from the ray to the Gaussian distribution, $S$ as the Gaussian scaling, and $R$ as the rotation matrix in the world space. Additionally, given $\Sigma=RSS^{\top}R^{\top}$ as the 3D covariance matrix of the Gaussian in the world space, we have $\mathrm{W}_{\mathrm{pca}}$ as the PCA whitening matrix satisfying:
(C.1) $\mathrm{W}_{\mathrm{pca}}$ $=\Sigma^{-\frac{1}{2}}=S^{-1}R^{\top},$
$\text{where}\quad S^{-1}$ $=\text{diag}\left(\frac{1}{\mathbf{s}}\right).$
As illustrated in Figure 4, we can apply the chain rule to find derivatives w.r.t. the Gaussian parameters, e.g, scaling vector $\mathbf{s}$, rotation quaternion $\mathbf{q}$, and mean $\boldsymbol{\mu}$:
(C.2) $\frac{\partial T}{\partial\mathbf{s}}$ $=\frac{\partial T}{\partial\kappa}\frac{\partial\kappa}{\partial\mathrm{W}_{\mathrm{pca}}}\frac{\partial\mathrm{W}_{\mathrm{pca}}}{\partial\mathbf{s}},$
(C.3) $\frac{\partial T}{\partial\mathbf{q}}$ $=\frac{\partial T}{\partial\kappa}\frac{\partial\kappa}{\partial\mathrm{W}_{\mathrm{pca}}}\frac{\partial\mathrm{W}_{\mathrm{pca}}}{\partial\mathbf{q}},$
(C.4) $\frac{\partial T}{\partial\boldsymbol{\mu}}$ $=\frac{\partial T}{\partial\kappa}\frac{\partial\kappa}{\partial\boldsymbol{\mu}}.$
C.2. Two-Stage Propagation
To reduce computational overhead in shared propagation, we design our gradient computation following the structured two-stage approach of 3DGS (Kerbl et al., 2023), comprising RenderCUDA (RC) and PreprocessCUDA (PC). The former performs ray-wise accumulation, efficiently accelerated by our Particle Bounding Frustum (PBF) association (See section 3.3), while the latter operates with Gaussian-wise complexity, leveraging view frustum culling for additional speedup.
According to Eq. 7, we can respectively write the gradient w.r.t. opacity coefficient and the shared partial derivative for Gaussian parameters in RC as:
(C.5) $\frac{\partial T}{\partial\sigma}\bigg{|}_{\texttt{RC}}=\alpha,\quad\frac{\partial T}{\partial\kappa}\bigg{|}_{\texttt{RC}}=-\frac{T}{2}.$
The remaining gradients to the Gaussian parameters can then be back-propagated from the squared distance, e.g., $\frac{\partial\kappa}{\partial\mathbf{s}},\frac{\partial\kappa}{\partial\mathbf{q}},\frac{\partial\kappa}{\partial\boldsymbol{\mu}}$.
C.3. Gradients for Canonical Ray
We then backpropagate gradients to the canonical ray parameters: the origin $\mathbf{o}_{u}$, direction $\mathbf{d}_{u}$, and moment $\mathbf{m}_{u}=\mathbf{o}_{u}\times\mathbf{d}_{u}$, as defined in Eq. 8 in our paper. These quantities directly influence the Mahalanobis distance in Eq. 9 and are critical for capturing canonical ray geometry.
(C.6) $\frac{\partial\kappa}{\partial\mathbf{m}_{u}}\bigg{|}_{\texttt{RC}}$ $=\left(\frac{2}{\mathbf{d}_{u}^{\top}\mathbf{d}_{u}}\right)\mathbf{m}_{u},$
(C.7) $\frac{\partial\kappa}{\partial\mathbf{o}_{u}}\bigg{|}_{\texttt{RC}}$ $=\frac{\partial\kappa}{\partial\mathbf{m}_{u}}\times\mathbf{d}_{u},$
(C.8) $\frac{\partial\kappa}{\partial\mathbf{d}_{u}}\bigg{|}_{\texttt{RC}}$ $=-\left(\frac{2\kappa}{\mathbf{d}_{u}^{\top}\mathbf{d}_{u}}\right)\mathbf{d}_{u}+\mathbf{o}_{u}\times\frac{\partial\kappa}{\partial\mathbf{m}_{u}}.$
C.4. Gradients for Whitening Matrix
Recall that $\mathbf{d}_{u}=\mathrm{W}_{\mathrm{pca}}\mathbf{d}$ and $\mathbf{o}_{u}=\mathrm{W}_{\mathrm{pca}}(\mathbf{o}-\boldsymbol{\mu})$, where $\mathbf{o}=-R_{c}^{\top}\mathbf{t}_{c}$ denotes the world space optical center, and $\mathbf{d}$ denotes the ray direction. Once the gradients w.r.t. the ray parameters are obtained, we further propagate them to the PCA whitening matrix $\mathrm{W}_{\mathrm{pca}}$ via outer products with the world-space vectors:
(C.9) RenderCUDA: $\quad\frac{\partial\kappa}{\partial\mathrm{W}_{\mathrm{pca}}}\bigg{|}_{\texttt{RC}}=\frac{\partial\kappa}{\partial\mathbf{d}_{u}}\mathbf{d}^{\top};$
PreprocessCUDA: $\quad\frac{\partial\kappa}{\partial\mathrm{W}_{\mathrm{pca}}}\bigg{|}_{\texttt{PC}}=\frac{\partial\kappa}{\partial\mathbf{o}_{u}}(\mathbf{o}-\boldsymbol{\mu})^{\top};$
Overall: $\quad\frac{\partial\kappa}{\partial\mathrm{W}_{\mathrm{pca}}}=\frac{\partial\kappa}{\partial\mathrm{W}_{\mathrm{pca}}}\bigg{|}_{\texttt{RC}}+\frac{\partial\kappa}{\partial\mathrm{W}_{\mathrm{pca}}}\bigg{|}_{\texttt{PC}}.$
Here, the gradient backpropagated through $\mathbf{d}_{u}$ is accumulated ray-wise in RenderCUDA, while the gradient through $\mathbf{o}_{u}$ is computed Gaussian-wise in PreprocessCUDA. These two components are then summed to yield the complete gradient w.r.t. the PCA transformation matrix $\mathrm{W}_{\mathrm{pca}}$.
C.5. Gradients for Scaling and Rotation
Thus, for scaling gradients $\frac{\partial\kappa}{\partial\mathbf{s}}=\frac{\partial\kappa}{\partial\mathrm{W}_{\mathrm{pca}}}\frac{\partial\mathrm{W}_{\mathrm{pca}}}{\partial\mathbf{s}}$, we have:
(C.10) $\frac{\partial\mathrm{W}_{\mathrm{pca}}^{i,j}}{\partial s_{k}}\bigg{|}_{\texttt{PC}}$ =$=\begin{Bmatrix}-\frac{1}{s_{k}^{2}}R_{j,k}&\mathrm{if}\ i=k\\ 0&\mathrm{otherwise}\end{Bmatrix}.$
To derive gradients for rotation $\frac{\partial\kappa}{\partial\mathbf{q}}=\frac{\partial\kappa}{\partial\mathrm{W}_{\mathrm{pca}}}\frac{\partial\mathrm{W}_{\mathrm{pca}}}{\partial\mathbf{q}}$, given the quaternion form $\mathbf{q}=q_{r}+q_{i}\mathbf{i}+q_{j}\mathbf{j}+q_{k}\mathbf{k}$, we can write the whitening matrix as:
(C.11) $\mathrm{W}_{\mathrm{pca}}=2\begin{pmatrix}\frac{1}{s_{1}}(\frac{1}{2}-q_{j}^{2}-q_{k}^{2})&\frac{1}{s_{1}}(q_{i}q_{j}+q_{r}q_{k})&\frac{1}{s_{1}}(q_{i}q_{k}-q_{r}q_{j})\\ \frac{1}{s_{2}}(q_{i}q_{j}-q_{r}q_{k})&\frac{1}{s_{2}}(\frac{1}{2}-q_{i}^{2}-q_{k}^{2})&\frac{1}{s_{2}}(q_{j}q_{k}+q_{r}q_{i})\\ \frac{1}{s_{3}}(q_{i}q_{k}+q_{r}q_{j})&\frac{1}{s_{3}}(q_{j}q_{k}-q_{r}q_{i})&\frac{1}{s_{3}}(\frac{1}{2}-q_{i}^{2}-q_{j}^{2})\end{pmatrix}.$
As a result, we derive the following end gradients for the quaternion $\mathbf{q}$ as:
(C.14) $\frac{\partial\mathrm{W}_{\mathrm{pca}}}{\partial q_{r}}\bigg{|}_{\texttt{PC}}=2\begin{pmatrix}0&\frac{q_{k}}{s_{1}}&-\frac{q_{j}}{s_{1}}\\ -\frac{q_{k}}{s_{2}}&0&\frac{q_{i}}{s_{2}}\\ \frac{q_{j}}{s_{3}}&-\frac{q_{i}}{s_{3}}&0\end{pmatrix},\hfil\hskip 0.0pt&\frac{\partial\mathrm{W}_{\mathrm{pca}}}{\partial q_{i}}\bigg{|}_{\texttt{PC}}=2\begin{pmatrix}0&\frac{q_{j}}{s_{1}}&\frac{q_{k}}{s_{1}}\\ \frac{q_{j}}{s_{2}}&-\frac{2q_{i}}{s_{2}}&\frac{q_{r}}{s_{2}}\\ \frac{q_{k}}{s_{3}}&-\frac{q_{r}}{s_{3}}&-\frac{2q_{i}}{s_{3}}\end{pmatrix},\\ \frac{\partial\mathrm{W}_{\mathrm{pca}}}{\partial q_{j}}\bigg{|}_{\texttt{PC}}=2\begin{pmatrix}-\frac{2q_{j}}{s_{1}}&\frac{q_{i}}{s_{1}}&-\frac{q_{r}}{s_{1}}\\ \frac{q_{i}}{s_{2}}&0&\frac{q_{k}}{s_{2}}\\ \frac{q_{r}}{s_{3}}&\frac{q_{k}}{s_{3}}&-\frac{2q_{j}}{s_{3}}\end{pmatrix},\hfil\hskip 0.0pt&\frac{\partial\mathrm{W}_{\mathrm{pca}}}{\partial q_{k}}\bigg{|}_{\texttt{PC}}=2\begin{pmatrix}-\frac{2q_{k}}{s_{1}}&\frac{q_{r}}{s_{1}}&\frac{q_{i}}{s_{1}}\\ -\frac{q_{r}}{s_{2}}&-\frac{2q_{k}}{s_{2}}&\frac{q_{j}}{s_{2}}\\ \frac{q_{i}}{s_{3}}&\frac{q_{j}}{s_{3}}&0\end{pmatrix}.\end{array}$
C.6. Gradients for 3D Mean
Meanwhile, we compute end gradients w.r.t. the Gaussian 3D mean $\boldsymbol{\mu}$ as:
(C.15) $\frac{\partial\kappa}{\partial\boldsymbol{\mu}}\bigg{|}_{\texttt{PC}}=-\mathrm{W}_{\mathrm{pca}}^{\top}\frac{\partial\kappa}{\partial\mathbf{o}_{u}}.$
Appendix D High-level Algorithmic Summary
We provide a high-level overview of our volumetric rendering pipeline (See Alg. 1 and Alg. 2). For clarity, we follow the structure of 3DGS (Kerbl et al., 2023) to illustrate differences: newly introduced modules or replaced ones are highlighted in green, and significantly modified ones in yellow.
Function 3DGEER($\boldsymbol{\mu},\mathbf{s},\mathbf{q},c,\sigma,\mathbf{r}_{c},V$):
     $\mathrm{CullGaussian}(\boldsymbol{\mu},V)$triangleright Frustum Culling
     $F\leftarrow\mathrm{InitCSFs}(w,h,K)$triangleright Partition
     $\Sigma_{c},\boldsymbol{\mu}_{c}\leftarrow\mathrm{ViewspaceGaussians}(\mathbf{s},\mathbf{q},V)$triangleright Transform
     $F_{ID}\leftarrow\mathrm{PBF\_Intersect}(\Sigma_{c},\boldsymbol{\mu},F)$triangleright Sub-frustum Asso.
     $\mathrm{W}_{\mathrm{pca}}\leftarrow\mathrm{PCA\_Whitening}(\mathbf{s},\mathbf{q})$triangleright Transform
     $\mathcal{G}_{ID},F_{K}\leftarrow\mathrm{DuplicateWithKeys}(\mathrm{W}_{\mathrm{pca}},F_{ID})$
      $\triangleright$ Indices (Gaussian) & Keys (FrustumID+Depth)
     $\mathrm{SortByKeys}(\mathcal{G}_{ID},F_{K})$triangleright Globally Depth Sort
     $R\leftarrow\mathrm{IdentifyFrustumRanges}(F,F_{K})$
     $I\leftarrow\mathbf{0}$triangleright Init Canvas (w/ Rays)
     for $\mathbf{all}\ \mathrm{CSFs}\ f\leftarrow I$ do
           for $\mathbf{all}\ \mathrm{Rays}\ i\leftarrow f$ do
                $r\leftarrow\mathrm{GetFrustumRange}(R,f)$
               $\mathbf{m}_{u},\mathbf{d}_{u}\leftarrow\mathrm{Cano\_Ray}(\mathrm{W}_{\mathrm{pca}},\boldsymbol{\mu},i)$triangleright Transform
               $\mathrm{D}_{\boldsymbol{\mu},\Sigma}\leftarrow\mathrm{Dist}(\mathbf{m}_{u},\mathbf{d}_{u})$triangleright Mah-distance
               $I[i]\leftarrow\mathrm{BlendInOrder}(i,\mathcal{G}_{ID},r,F_{K},\mathrm{D}_{\boldsymbol{\mu},\Sigma},c,\sigma)$
                $\triangleright$ Alpha Blending
                end for
                end for
               return $I$
ALGORITHM 1 GPU Software Asso. & Render.
$\triangleright$ Init Gaussian Attributes
$\boldsymbol{\mu}\leftarrow\mathrm{SfM\ Points}$triangleright Position
$\mathbf{s},\mathbf{q}\leftarrow\mathrm{InitAttributes}()$triangleright Scaling, Rotation
$c,\ \sigma\leftarrow\mathrm{InitCoefficients}()$triangleright Color, Opacity
while $\mathrm{not\ converged}$ do
       $\mathbf{r}_{c},\hat{C}\leftarrow\mathrm{BEAP}(w,h,K)$triangleright View Rays / Colors
       $\triangleright$ Asso. & Render
      $C\leftarrow\mathrm{3DGEER}(\boldsymbol{\mu},\mathbf{s},\mathbf{q},c,\sigma,\mathbf{r}_{c},V)$triangleright See Alg. 1
      $\mathcal{L}\leftarrow\mathrm{Loss}(C,\hat{C})$triangleright Loss
      $\boldsymbol{\mu},\mathbf{s},\mathbf{q},c,\sigma\leftarrow\mathrm{Adam}(\nabla\mathcal{L})$triangleright Backprop & Step
      if $\mathrm{IsRefinementIteration}(i)$ then
             $\boldsymbol{\mu},\mathbf{s},\mathbf{q},c,\sigma\leftarrow\mathrm{Optimize}(\boldsymbol{\mu},\mathbf{s},\mathbf{q},c,\sigma)$triangleright (Kerbl et al., 2023)
       end if
end while
$w,h$: width and height of the training images
$K,V$: intrinsics and extrinsic meta info.
ALGORITHM 2 3DGEER Initialization and Training Framework
$w,h$: width and height of the training images
$K,V$: intrinsics and extrinsic meta info.
Appendix E Additional Details in Association
E.1. Particle Bounding Frustum (PBF) Solver
Recall Eq. 12, we define $P=(VH)^{\top}\in\mathbb{R}^{4\times 4}$ as the plane transformation matrix from the view space to the canonical space. Take the unknown bounds for $c=\tan\theta_{1,2}$ as example, we have:
(E.1) $\mathbf{g}_{u}=-P_{0}+cP_{2}$ $=\begin{bmatrix}-P_{00}+cP_{02}\\ -P_{10}+cP_{12}\\ -P_{20}+cP_{22}\\ -P_{30}+cP_{32}\end{bmatrix},$
where $P_{i}$ denotes $i$-th row of the matrix, expressed as a $4\times 1$ column vector, and $P_{ij}$ further denotes $j$-th element. Recall Eq. 13, assuming we are computing the bounding box of the $1$-standard-deviation contour ellipsoid of the Gaussian, i.e., $\lambda=1$, then we have the following quadratic equation:
(E.2) $\left(\left(\mathbf{f}\circ P_{2}^{\top}\right)\cdot P_{2}\right)c^{2}-2\left(\left(\mathbf{f}\circ P_{0}^{\top}\right)\cdot P_{2}\right)c+\left(\mathbf{f}\circ P_{0}^{\top}\right)\cdot P_{0}=0,$
where $\mathbf{f}=(1,1,1,-1)$ is a $1\times 4$ row vector. Here we denote the equation as:
(E.3) $\mathcal{T}_{22}c^{2}-2\mathcal{T}_{02}c+\mathcal{T}_{00}=0,$
where $\mathcal{T}$ is a $3\times 3$ upper triangular matrix defined with $\mathcal{T}_{ij}=\left(f\circ P_{i}^{\top}\right)\cdot P_{j}$.
Similarly, for the bounds of $\phi$, we have $c=\tan\phi_{1,2}$ satisfying:
(E.4) $\mathcal{T}_{22}c^{2}-2\mathcal{T}_{12}c+\mathcal{T}_{11}=0.$
Interestingly, given a camera-view configuration $[R_{c}\mid\mathbf{t}_{c}]$, we find that the parameter matrix $\mathcal{T}$ can be efficiently computed from the Gaussian’s view-space covariance $\Sigma_{c}=R_{c}RSS^{\top}R^{\top}R_{c}^{\top}$ and 3D mean $\boldsymbol{\mu}_{c}=R_{c}\boldsymbol{\mu}+\mathbf{t}_{c}$:
(E.5) $\mathcal{T}_{ij}=\lambda^{2}\Sigma_{c}^{i,j}-\left(\boldsymbol{\mu}_{c}\boldsymbol{\mu}_{c}^{\top}\right)^{i,j}.$
Thus, we can compute the angular bounds (defined by their tangent-space center and extent) from:
(E.6) $\left(\frac{\tan\theta_{1}+\tan\theta_{2}}{2},\frac{\tan\phi_{1}+\tan\phi_{2}}{2}\right)=\left(\frac{\mathcal{T}_{02}}{\mathcal{T}_{22}},\frac{\mathcal{T}_{12}}{\mathcal{T}_{22}}\right),$
$\|\tan\theta_{1}-\tan\theta_{2}\|=\frac{2\sqrt{\mathcal{T}_{02}^{2}-\mathcal{T}_{22}\mathcal{T}_{00}}}{|\mathcal{T}_{22}|},$
$\|\tan\phi_{1}-\tan\phi_{2}\|=\frac{2\sqrt{\mathcal{T}_{12}^{2}-\mathcal{T}_{22}\mathcal{T}_{11}}}{|\mathcal{T}_{22}|}.$
If the discriminant of either quadratic is negative—indicating that the camera’s optical center lies inside the Gaussian ellipsoid and the angular bounds are undefined—we conservatively clamp the frustum angles using the camera’s maximum field of view (FoV).
Note that this result is directly computed from the 3D view-space true covariance, in contrast to methods like EWA and UT, which rely on intermediate conic approximations in screen space. Our PBF formulation offers both exactness and efficiency.
E.2. Details in Computing CSF-PBF Intersection
Figure E.1. Mirror Transformation of PBF Angular Bounds. Four candidate mirrored angles $\theta_{1,\xi},\theta^{\prime}_{1,\xi},\theta_{2,\xi},\theta^{\prime}_{2,\xi}$ are generated from a given pair of bounds. Given a Gaussian centered at $\boldsymbol{\mu}_{c}$, the two closest to the corresponding Gaussian’s mirrored angle $\theta_{\mu,\xi}$ are used as PBF bounds for robust intersection with mirror-transformed CSF partitions.
We uniformly sample spherical angles within the maximum field of view (FoV). The resulting $x$, $y$, and $z$ coordinates—i.e., the ray direction vectors $\mathbf{d}$—are projected onto a unit sphere and then transformed into pixel space for inverse color interpolation (see Appendix F). These sampled rays are then grouped according to their corresponding CSFs for subsequent Gaussian association.
In most scenes using a pinhole camera with a limited FoV, the angular bounds in PBF typically lie within $[-\frac{\pi}{2},\frac{\pi}{2}]$. In such cases, since the $\tan$ function is monotonically increasing over this interval, the intersections between PBF and CSFs can be computed by directly comparing the $\tan$ values of their respective angular bounds. However, in wide FoV scenarios, even when the incidence angle formed by the Gaussian centers remains within $[-\frac{\pi}{2},\frac{\pi}{2}]$, the angular bounds used in PBF may still fall outside the monotonic region of the $\tan$ function. In these cases, directly comparing $\tan$ values may yield incorrect PBF-CSF intersections, such as mistakenly matching angular bounds as $[\theta^{\prime}_{2},\theta_{1}]$ (see Figure E.1 for an example).
To resolve the correct intersections, we introduce a mirror transformation applied to the tangent values of both PBF angular bounds and CSF partitions. Inspired by the CMEI camera model (Mei and Rives, 2007), we shift the optical center using a mirror parameter $\xi$, leading to the following transformation:
(E.7) $\tan\theta_{\xi}=\frac{\tan\theta}{1+\xi\frac{z}{\|z\|}\sqrt{1+\tan^{2}\theta}}=\frac{\sin\theta}{\cos\theta+\xi},$
where $\theta_{\xi}$ represents the mirror-transformed angle of $\theta$ on the $xz$-plane, using a mirror parameter $\xi$. We set $\xi=1$ throughout all experiments.
Given a pair of tangent value $[\tan\theta_{1},\tan\theta_{2}]$ from our PBF solver (See Equation. E.6), this transformation yields four candidate mirror angles, as illustrated in Figure E.1. The Gaussian center, in contrast, corresponds to a unique mirror-transformed angle $\theta_{\mu,\xi}$. We exploit this property by selecting the two candidate angles closest to $\theta_{\mu,\xi}$. These two then define the effective lower and upper bounds for PBF, i.e., $[\theta_{1,\xi},\theta_{2,\xi}]$, which are subsequently compared with CSF partitions to compute the correct intersections.
This trick ensures robust and consistent angular comparisons, even beyond the monotonic domain of the $\tan$ function. As reported in our timing analysis (Table 6), the pre-processing pipeline—including the mirror-angle-based PBF-CSF intersection computation and the canonical transformation for all Gaussians—takes less than 0.13 ms in total. This confirms the efficiency of our approach, even when applied to thousands of Gaussians across wide FoV scenarios.
Appendix F BEAP Image Inverse Interpolation
To prepare training data, we sample rays to comprehensively cover the original image FoVs and use them as input to our CUDA renderer. Simultaneously, we interpolate the ground truth color for each input ray to ensure accurate and uniform supervision during training.
The projection of sampled rays onto observed images begins with transforming the incidence angle projections $\theta$ and $\phi$ into normalized ray direction vectors $\mathbf{d}=(x,y,z)^{\top}$, expressed as:
(F.1) $x$ $=\frac{\sin\theta\cos\phi}{\sqrt{\sin^{2}\theta\cos^{2}\phi+\cos^{2}\theta\sin^{2}\phi+\cos^{2}\theta\cos^{2}\phi}},$
(F.2) $y$ $=\frac{\cos\theta\sin\phi}{\sqrt{\sin^{2}\theta\cos^{2}\phi+\cos^{2}\theta\sin^{2}\phi+\cos^{2}\theta\cos^{2}\phi}},$
(F.3) $z$ $=\frac{\cos\theta\cos\phi}{\sqrt{\sin^{2}\theta\cos^{2}\phi+\cos^{2}\theta\sin^{2}\phi+\cos^{2}\theta\cos^{2}\phi}}.$
Several camera models are provided to subsequently project the normalized rays $(x,y,z)^{\top}$ into pixel space $(x_{p},y_{p},1)^{\top}$.
Pinhole Camera Model
For pinhole camera models, we first have $\frac{x}{z}=\tan\theta$ and $\frac{y}{z}=\tan\phi$. Thus, we have:
(F.4) $\begin{bmatrix}x_{p}\\ y_{p}\\ 1\end{bmatrix}=\begin{bmatrix}f_{x}&0&c_{x}\\ 0&f_{y}&c_{y}\\ 0&0&1\end{bmatrix}\begin{bmatrix}\tan\theta\\ \tan\phi\\ 1\end{bmatrix}.$
KB Fisheye Model
We have the incidence angle $\varphi$ derived from its projections:
(F.5) $r=\sqrt{\tan^{2}\theta+\tan^{2}\phi}\quad\text{and}\quad\varphi=\arctan(r).$
And we have the fisheye distorted angle:
(F.6) $\varphi_{d}=\varphi(1+k_{1}\varphi^{2}+k_{2}\varphi^{4}+k_{3}\varphi^{6}+k_{4}\varphi^{8}).$
As a result, the rays are projected onto pixel space, allowing for inverse color interpolation from:
(F.7) $\begin{bmatrix}x_{p}\\ y_{p}\\ 1\end{bmatrix}=\begin{bmatrix}\frac{f_{x}\varphi_{d}}{r}&0&c_{x}\\ 0&\frac{f_{y}\varphi_{d}}{r}&c_{y}\\ 0&0&1\end{bmatrix}\begin{bmatrix}\tan\theta\\ \tan\phi\\ 1\end{bmatrix}.$
Appendix G Additional Implementation Details
Central and peripheral regions in ScanNet++ evaluation. The rays sampled in the central part follows the range:
$\theta\in[-\frac{\mathrm{w}}{2f_{x}},\frac{\mathrm{w}}{2f_{x}}],\phi\in[-\frac{\mathrm{h}}{2f_{y}},\frac{\mathrm{h}}{2f_{y}}].$
And the Peripheral part takes the rays sampled in remaining angle ranges. Here $\mathrm{w},\mathrm{h}$ indicate the width and height of the image and $f_{x},f_{y}$ are the focal lengths defined in the original KB camera model.
Cross-camera generalization evaluation on ZipNeRF. ZipNeRF provides two sets of data: fisheye raw images and their corresponding undistorted frames, which are generated by applying distortion correction and cropping the central regions. Although the frames are aligned on a per-frame basis, camera poses and point clouds were independently reconstructed using COLMAP (Schönberger and Frahm, 2016) for each set. Due to COLMAP’s global optimization independently refining camera trajectories, applying a single rigid transformation to align the two pose sets is insufficient and may introduce alignment errors. To ensure reliable cross-camera evaluation without introducing errors from structure-from-motion (SfM), we always use the same set of test-time camera poses as those used during training. In cross-camera experiments, we vary only the intrinsic parameters, while keeping the extrinsic poses fixed. This ensures fair comparisons without entangling pose inconsistencies arising from preprocessing.
Appendix H Additional Results
See Next Page for Full Experiment Results.
Table H.1. ScanNet++ Quantitative Results by Sequence (Full FoV). The overall results are shown in Tab. 2.
0a5c013435 2a1a3afad9 1f7cbbdde1 4ef75031e3 1d003b07bd
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
EVER 28.24 0.916 0.164 27.80 0.899 0.169 32.60 0.948 0.131 29.77 0.924 0.213 28.96 0.936 0.156
FisheyeGS 24.74 0.932 0.146 25.75 0.927 0.145 31.95 0.970 0.095 29.72 0.944 0.177 26.91 0.959 0.129
3DGUT 29.26 0.934 0.147 29.37 0.928 0.145 34.65 0.967 0.112 30.78 0.941 0.196 29.12 0.948 0.151
Ours 29.18 0.940 0.141 29.78 0.935 0.125 34.94 0.974 0.084 31.17 0.946 0.168 32.43 0.968 0.114
Table H.2. ZipNeRF Quantitative Results by Sequence. Alameda, Berlin, London, and NYC in order. The overall results are shown in Tab. 3.
Train Data $1/8$ FE $1/8$ FE $1/8$ FE $1/4$ PH $1/4$ PH $1/4$ PH
Test Data $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
FisheyeGS 20.74 23.64 0.820 0.802 0.223 0.277 17.39 23.74 0.736 0.839 0.265 0.240
EVER 22.46 23.06 0.852 0.777 0.144 0.274 20.06 22.83 0.789 0.794 0.213 0.250
3DGUT 21.78 22.39 0.841 0.718 0.193 0.368 17.03 22.58 0.588 0.765 0.329 0.311
Ours 24.60 25.00 0.857 0.839 0.154 0.241 21.49 25.21 0.792 0.807 0.232 0.277
Train Data $1/8$ FE $1/8$ FE $1/8$ FE $1/4$ PH $1/4$ PH $1/4$ PH
Test Data $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
FisheyeGS 24.25 28.25 0.895 0.924 0.206 0.209 20.08 27.86 0.848 0.929 0.233 0.194
EVER 25.26 26.97 0.912 0.907 0.155 0.221 23.18 27.72 0.877 0.913 0.196 0.212
3DGUT 25.37 27.02 0.904 0.876 0.197 0.288 18.57 27.43 0.730 0.898 0.297 0.249
Ours 25.51 29.15 0.916 0.925 0.166 0.194 23.58 29.51 0.878 0.915 0.219 0.228
Train Data $1/8$ FE $1/8$ FE $1/8$ FE $1/4$ PH $1/4$ PH $1/4$ PH
Test Data $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
FisheyeGS 23.57 26.65 0.829 0.861 0.253 0.254 20.83 27.33 0.768 0.883 0.288 0.229
EVER 24.59 26.83 0.823 0.859 0.218 0.223 23.14 26.98 0.793 0.860 0.268 0.235
3DGUT 25.61 26.25 0.854 0.803 0.221 0.336 19.82 26.41 0.669 0.833 0.313 0.291
Ours 26.60 27.84 0.895 0.882 0.142 0.230 23.95 27.15 0.840 0.846 0.217 0.289
Train Data $1/8$ FE $1/8$ FE $1/8$ FE $1/4$ PH $1/4$ PH $1/4$ PH
Test Data $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH $1/8$ FE $1/4$ PH
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
FisheyeGS 24.15 27.21 0.890 0.885 0.162 0.215 19.41 27.50 0.811 0.907 0.203 0.183
EVER 28.39 27.71 0.931 0.860 0.095 0.228 25.22 28.27 0.879 0.883 0.154 0.192
3DGUT 26.31 26.69 0.917 0.817 0.121 0.302 19.00 27.42 0.659 0.866 0.263 0.228
Ours 28.26 28.49 0.922 0.906 0.097 0.190 24.52 28.55 0.874 0.885 0.168 0.220
Table H.3. MipNeRF360 Quantitative Results by Sequence. The overall results are shown in Tab. 4.
Method Metric Bicycle Bonsai Counter Garden Kitchen Stump Flowers Room Treehill
[…]
Appendix I EWA Splatting Preliminaries - Local Approximation Error
From camera coordinates to ray coordinates As the coordinate definitions in EWA Splatting (Zwicker et al., 2001), we have the following mapping function from the camera (view) space points $\mathbf{t}$ to the ray space points $\mathbf{x}$ for projective transformation in pinhole camera models, i.e., $\tau:\ \mathbf{t}\xrightarrow{\sim}\mathbf{x}$:
(I.1) $\mathbf{x}$ $=\begin{bmatrix}x_{0}\\ x_{1}\\ x_{2}\end{bmatrix}=\tau(\mathbf{t})=\begin{bmatrix}t_{0}/t_{2}\\ t_{1}/t_{2}\\ \left\|(t_{0},t_{1},t_{2})^{\top}\right\|\end{bmatrix},$
(I.2) $\mathbf{t}$ $=\begin{bmatrix}t_{0}\\ t_{1}\\ t_{2}\end{bmatrix}=\tau^{-1}(\mathbf{x})=\begin{bmatrix}x_{0}/l\cdot x_{2}\\ x_{1}/l\cdot x_{2}\\ 1/l\cdot x_{2}\end{bmatrix},$
(I.3) $,\ \mathrm{where}\ l$ $=\left\|(x_{0},x_{1},1)^{\top}\right\|.$
These mappings are not affine, which requires computationally expensive (Westover, 1990; Mao et al., 1995) or locally approximated solutions (Zwicker et al., 2001; Kerbl et al., 2023) to calculate integral in the ray space.
Local affine approximation EWA splatting by Zwicker et al. (2001) introduced local affine approximation, denoted as $\tau_{k}$, to keep the linear effects and the $k$-indexed reconstruction kernels as Gaussians in 2D screen space. It is defined by the first two terms of the Taylor expansion of $\tau$ at the camera space point $\mathbf{t}_{k}$, then the ray space point $\mathbf{x}$ near around $\mathbf{x}_{k}$ can be estimated by:
(I.4) $\mathbf{x}$ $=\tau(\mathbf{t})=\mathbf{x}_{k}+\mathbf{J}_{\tau}^{\mathbf{t}_{k}}(\mathbf{t}-\mathbf{t}_{k}),$
where $\mathbf{J}_{\tau}^{\mathbf{t}_{k}}$ is the Jacobian matrix given by the partial derivatives of $\tau$ at the point $\mathbf{t}_{k}$. For the $k$-indexed 3D Gaussian $\mathcal{G}_{\Sigma,\boldsymbol{\mu}}$ in a certain camera coordinate, using EWA local affine transformation, the approximated density function can be expressed as a 3D Gaussian $\mathcal{G}_{\Sigma_{k},\mathbf{x}_{k}}$ in the ray coordinate. For $\forall\mathbf{x}$ around $\mathbf{x}_{k}$, we have:
(I.5) $\mathcal{G}_{\Sigma_{k},\mathbf{x}_{k}}(\mathbf{x})$ $=\frac{1}{(2\pi)^{\frac{3}{2}}|\Sigma_{k}|^{\frac{1}{2}}}\exp\left({-\frac{1}{2}(\mathbf{x}-\mathbf{x}_{k})^{\top}\Sigma_{k}^{-1}(\mathbf{x}-\mathbf{x}_{k})}\right),$
(I.6) $\mathrm{where}\ \Sigma_{k}$ $=\mathbf{J}_{\tau}^{\mathbf{t}_{k}}\Sigma\left(\mathbf{J}_{\tau}^{\mathbf{t}_{k}}\right)^{\top}.$
Integrating a 3D Gaussian $\mathcal{G}_{\Sigma_{k},\mathbf{x}_{k}}^{(3)}$ along one coordinate axis yields a 2D Gaussian $\mathcal{G}^{(2)}_{\hat{\Sigma_{k}},\hat{\mathbf{x}}_{k}}$ , hence:
[…]
where $\hat{\mathbf{x}}=\begin{bmatrix}x_{0}\\ x_{1}\end{bmatrix}$, and the mean $\hat{\mathbf{x}}_{k}=\begin{bmatrix}x_{k,0}\\ x_{k,1}\end{bmatrix}$, and the $2\times 2$ variance matrix $\hat{\Sigma}_{k}$ is easily obtained from the $3\times 3$ matrix $\Sigma_{k}$ by skipping the third row and column.
Splatting approximation error Although we can not expect the density function in ray space still obey the Gaussian distribution when we apply general projective transformation to the original view space, exact density can be derived through substitution via Eq. I.2:
(I.8) $\mathcal{G}_{\Sigma,\boldsymbol{\mu}}(\mathbf{t})$ $=\mathcal{G}_{\Sigma,\boldsymbol{\mu}}\left(\tau^{-1}(\mathbf{x})\right)$
$=\frac{1}{(2\pi)^{\frac{3}{2}}|\Sigma_{k}|^{\frac{1}{2}}}\exp\left({-\frac{1}{2}\left(\tau^{-1}(\mathbf{x})-\mathbf{t}_{k}\right)^{\top}\Sigma_{k}^{-1}\left(\tau^{-1}(\mathbf{x})-\mathbf{t}_{k}\right)}\right),$
which means that the approximation error can be estimated through Monte Carlo integration of the following equation:
[…]
Note that $x_{2}$ represents the dimension along the ray direction in the ray coordinate.